# SPECIAL RELATIVITY

### v9

@@ 14.2 An Illustrative Example
In the next chapter we present a critical reexamination of existing synchrotron radiation theory. But before the discussion of this topic it would be well to illustrate error in standard coupling fields and particles in accelerator and plasma physics by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties. This illustrative example is mainly addressed to readers with limiting knowledge of accelerator and synchrotron radiation physics. Fortunately, the error in standard coupling fields and particles can be explained in a very simple way.
14.2.1 The Velocity of an Electron Accelerated by the Kicker Field
@@
An electron kicker setup is a practical case of study for illustrating the difference between covariant and non-covariant trajectories. Let us consider the simple case when an ultrarelativistic electron moving with the velocity $v$ along $z$-axis in the lab frame is kicked by a weak dipole magnetic field directed along $x$-axis. We assume for simplicity that the kick angle is small compare with $1/\gamma$, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the relativistic factor. This means that we take the limit $\gamma\gg 1\gg\gamma v_{y}/v$. Let us start with non-covariant particle tracking calculations. The trajectory of the electron, which follows from the solution of the corrected Newton’s second law under the absolute time convention, does not include relativistic effects. Therefore, as usual for Newtonian kinematics, Galilean vectorial law of addition of velocities is actually used. Non-covariant particle dynamics shows that the electron direction changes after the kick, while the speed remains unvaried (Fig. 60). According to non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ is only capable of altering the direction of motion, but not the speed of the electron. After the kick, the beam velocity components are $(0,v_{y},v_{z})$, where $v_{z}=\sqrt{v^{2}-v_{y}^{2}}$. Taking the ultrarelativistic limit $v\simeq c$ and using the second order approximation we get $v_{z}=v[1-v_{y}^{2}/(2v^{2})]=v[1-v_{y}^{2}/(2c^{2})]$.
@@
Figure 60: A setup for illustration the difference between covariant and non-covariant trajectories. The motion of a relativistic electron accelerated by a kicker field. It is assumed for simplicity that $\gamma\gg 1\gg\gamma v_{y}/v$. According to the non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ is only capable altering the direction of motion, but not the speed of the electron.
@@
In contrast, covariant particle tracking, which is based on the use of Lorentz coordinates, yields different results for the velocity of the electron. Let us consider a composition of passive Lorentz transformations that track the motion of the relativistic electron accelerated by the kicker field. Let the $S$ be the lab frame of reference and $S^{\prime}$ a comoving frame with velocity $\vec{v}$ relative to $S$. Upstream of the kicker, the particle is at rest in the frame $S^{\prime}$. In order to have this, we impose that $S^{\prime}$ is connected to $S$ by the Lorentz boost $L(\vec{v})$, with $\vec{v}$ parallel to the $z$ axis, which transforms a given four vector event $X$ in a space-time into $X^{\prime}=L(\vec{v})X$. Let us analyze the particle evolution within $S^{\prime}$ frame. Our particle is at rest and the kicker is running towards it with velocity $-\vec{v}$. The moving magnetic field of the kicker produces an electric field orthogonal to it. When the kicker interacts with the particle in $S^{\prime}$ we thus deal with an electron moving in the combination of perpendicular electric and magnetic fields.
@@
First, we analyze textbook presentation of a composition of motions. Let $S"$ be a frame fixed with respect to the particle downstream the kicker. As is known, the non collinear Lorentz boosts does not commute. In our second order approximation we can neglect the difference between the $\gamma v_{y}/c$ and $\gamma_{z}v_{y}/c$, where $\gamma_{z}=1/\sqrt{1-v_{z}^{2}/c^{2}}$. Here $v_{z}=v(1-\theta_{k}^{2}/2)$ and $\theta_{k}=v_{y}/v=v_{y}/c$ in our (ultrarelativistic) case of interest. Therefore we can use a sequence of two commuting non-collinear Lorentz boosts linking $X^{\prime}$ in $S^{\prime}$ to $X^{\prime\prime}$ in $S^{\prime\prime}$ as $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})X^{\prime}=L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e^{\prime}}_{y}v^{\prime}_{y})X^{\prime}$ in order to discuss the beam motion in the frame $S^{\prime}$ after the kick. Here $\vec{e^{\prime}}_{y}$ and $\vec{e^{\prime}}_{z}$ are unit vectors directed, respectively, along the $x^{\prime}$ and $z^{\prime}$ axis. Note that as observed by an observer on $S^{\prime}$, the axes of the frame $S^{\prime\prime}$ are parallel to those of $S^{\prime}$. The relation $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)X$ presents a step-by-step change from $S$ to $S^{\prime}$ and then to $S"$. For the simple case of parallel velocities, the addition law is $L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)=L(\vec{e}_{z}v_{z})$. The resulting (passive) Lorentz boost composition can be represented as $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e}_{z}v_{z})X$ and textbooks state that the magnetic field, in the Lorentz coordinatization, is only capable altering the direction of motion, but not the speed of the electron.
@@
According to textbooks, in relativistic kinematics, the composition of the perpendicular velocities is studied analytically from the composition of Lorentz boosts. We have already discussed the law of composition of velocities in Section 13.5. It was shown that velocities addition in the Lorentz coordinatization is regulated by the Wigner rotation. In the ultrarelativistic approximation we find the simple result $v_{tot}=v_{z}=v[1-v_{y}^{2}/(2c^{2})]$, so that a Lorentz transformation leads to a rotation of the particle velocity $v_{z}$ of the angle $v_{y}/c$ (Fig. 61). It becomes clear that the contradiction is related to the Wigner rotation. In fact, as we already discussed in the previous chapter, the relation $dt^{\prime}=dt/\gamma$ cannot be used when we deal with relativistic composition of perpendicular velocities.
@@
It should be note, however, that there is another satisfactory way of the covariant electron tracking. Let us now return to our consideration of the motion of a relativistic electron accelerated by the kicker field and let us discuss the observation of an inertial lab observer without passing from one reference frame to another. We now consider the case when the electron in the lab frame is accelerated by the kicker field. In other words, we consider an active boost. In the case of an active boost we are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system. The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the electron was upstream the kicker. This synchronization convention preserves simultaneity and is actually based on the absolute time (or absolute simultaneity) convention. The usual Galileo rule of addition of velocities is used to determine the Galileo boosts tracking an electron, instant after instant, along its motion in a constant magnetic field.
@@
In the Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization. The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame. Let us analyze the resynchronization process of the lab distant clocks during the acceleration of the electron. This will allow us to demonstrate a direct relation between the decrease of the electron speed after the kick in Lorentz coordinates and the time dilation phenomenon. As we already remarked, the Lorentz coordinate system is only a mental construction: manipulations with non existing clocks are only needed for the application of the usual Maxwell’s equations for synchrotron radiation calculations.
@@
Suppose that upstream the kicker we pick a Lorentz coordinates in the lab frame. Then, an instant after entering the magnetic field, the electron velocity changes of the infinitesimal value $d\vec{v}$ along the $y$-axis. At this first step, Eq.(21) allows us to express the differential $d\vec{v}$ through the differential $dt$ in the Lorentz coordinate system assigned upstream the kicker. If clock synchronization is fixed, this is equivalent to the application of the absolute time convention. In order to keep Lorentz coordinates in the lab frame, as discussed before, we need to perform a clock resynchronization by introducing an infinitesimal time shift. The simplest case is when the kick angle $\theta_{k}$ is very small, and we evaluate transformations, working only up to the order $(\theta_{k}\gamma)^{2}$. The restriction to this order provides an essential simplicity of calculations in our case of interest for two reasons. First, relativistic correction to compositions of non-collinear velocity increments does not appear in this expansion order, but only in the order $(\gamma\theta_{k})^{3}$. Second, the time dilation appears in the highest order we use. Thus, Eq.(21) allows us to express the small velocity change $\Delta\vec{v}$ after the kick in the initial Lorentz coordinates system, and to perform clock resynchronization only downstream the kicker. Therefore, after the kick we can consider the composition of two Lorentz boosts along the perpendicular $x$ and $z$ directions. The first boost imparts the velocity $v\theta_{k}\vec{e}_{y}$ to the electron along the $y$-axis and the second boost imparts the additional velocity $-(v\theta_{k}^{2}/2)\vec{e}_{z}$ along the $z$ axis, while the restriction to second order assures that the boosts commute.
@@
Figure 61: The motion of a relativistic electron accelerated by the kicker field. In order to describe the kicker influence on the electron trajectory as Lorentz transformation, one needs to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of all clocks, from $t$ to $\gamma_{y}t$, with $\gamma_{y}=1+\theta_{k}^{2}/2$. It follows that the total electron speed in the Lorentz lab frame downstream the kicker decreases from $v$ to $v(1-\theta^{2}_{k}/2)$.
@@
In order to keep a Lorentz coordinates system in the lab frame after the kick, that is equivalent to describe the kicker influence on the electron trajectory as Lorentz transformation, we need to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of all clocks, from $t$ to $\gamma_{y}t$, with $\gamma_{y}\simeq 1+\theta_{k}^{2}/2$. It is immediately understood that the speed of electron downstream the kicker is no longer independent of the electron motion in the magnetic field (Fig. 61). No relativistic correction to the velocity component along the $y$-axis appears in the second order, but a correction of the longitudinal velocity component, changing $v_{z}$ to $v_{z}/\gamma_{y}$ with $v_{z}=v(1-\theta_{k}^{2}/2)$ and $v_{z}/\gamma_{y}=v(1-\theta_{k}^{2})$. It follows that the total electron speed in the lab frame, after clock resynchronization downstream the kicker, decreases from $v$ to $v(1-\theta_{k}^{2}/2)$. The time dilation does not come into the calculation of the velocity increment, but appears in the correction of the initial (relativistic) velocity $\vec{v}=v\vec{e}_{z}$. We have found that we get the same result whether we using passive Lorentz transformations or using active Lorentz transformations as it must be.
Note that we discuss particle tracking in the limit of a small kick angle $\gamma v_{y}/c\ll 1$. However, even in this simple case and for a single electron we are able to demonstrate the difference between non-covariant and covariant particle trajectories. The electron speed decreases from $v$ to $v(1-\theta_{k}^{2}/2)$. This result is at odds with the prediction from non-covariant particle tracking.
14.2.2 Discussion
We now want to point out that when we accelerate the electron in the lab frame by a kicker, the information about this acceleration is included into the covariant trajectory. The result of covariant approach clearly depends on the absolute value of the kick angle. The argument concerning the relativity motion in our case of interest cannot be applied. Not all is relative in special relativity, because any change of velocity has an absolute meaning. The ”absolute” acceleration means acceleration with respect to the fixed stars.
There are several points to be made about the above results. One interesting question is, where physically is the information about the electron acceleration? We answer this question in great detail in the Section 15.6. It is a dynamical line of arguments that explains this paradoxical situation with the trajectory of the accelerated electron. Dynamics, based on the field equations, is actually hidden in the language of relativistic kinematics. Without proof, we may state the results. The distribution of the electromagnetic fields from a fast moving charge is described by Ginzburg-Frank formula (see Appendix IV). We remark first that for a rapidly moving electron we have nearly equal transverse and mutually perpendicular electric and magnetic fields. These are indistinguishable from the fields of a beam of radiation. From microscopic viewpoint, when the electron passes through a kicker, its fields are perturbed, and now include information about acceleration.
Let us quickly look at some other cases of interest. Above we consider the simple case when kick angle is small. When the electron passes through the weak kicker there is no synchrotron radiation (to be more precise, in this case radiation is indistinguishable from the self-electromagnetic fields of the electron). We expect that an electron that passes through the weak kicker is still ”field-dressed”. Suppose the ultrarelativistic electron in the lab frame is kicked by dipole magnet field, but the kick angle is large compare with $1/\gamma$. In other words, we now consider an electron moving along a standard synchrotron radiation setup. At the exit of strong kicker we have a ”naked” (or ”field-free”) electron i.e. an electron that is not accompanied by virtual radiation fields. There is a process of formation of the ”field dressed” electron (i.e. formation of the fields from a fast moving charge) within the formation length from the very beginning of the straight section downstream of the strong kicker. The information about acceleration is included now into the synchrotron radiation. But the acceleration history (together with the self fields of the ultrarelativistic electron) is washed out during the kicking process, and now electron fields not include information about acceleration. That is to say, the information about acceleration is not included into the covariant trajectory downstream the strong kicker (we will discuss this subject further in the Section 14.5).
14.2.3 The Motion of an Electron Accelerated by a Kicker in a Bending Magnet

### v10

@@ 18.2 An Illustrative Example
It would be well to illustrate errors in standard coupling fields and particles in accelerator and plasma physics by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties. This illustrative example is mainly addressed to readers with limited knowledge of accelerator and synchrotron radiation physics. Fortunately, the error in standard coupling fields and particles can be explained in a very simple way.
@@
An electron kicker setup is a practical case of study for illustrating the difference between covariant and non-covariant trajectories. Let us consider the simple case when an ultrarelativistic electron moving with the velocity $v$ along the $z$-axis in the inertial (lab) frame is kicked by a weak dipole magnetic field directed along the $x$-axis. We assume for simplicity that the kick angle is small compared with $1/\gamma$, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the relativistic factor. This means that we take the limit $\gamma\gg 1\gg\gamma v_{y}/v$. Let us start with non-covariant particle tracking calculations. The trajectory of the electron, which follows from the solution of the corrected Newton’s second law under the absolute time convention, does not include relativistic effects. Therefore, as usual for Newtonian kinematics, the Galilean vectorial law of addition of velocities is used. Non-covariant particle dynamics shows that the electron direction changes after the kick, while the speed remains unvaried (Fig. 49). According to non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ is only capable of altering the direction of motion, but not the speed of the electron. After the kick, the beam velocity components are $(0,v_{y},v_{z})$, where $v_{z}=\sqrt{v^{2}-v_{y}^{2}}$. Taking the ultrarelativistic limit $v\simeq c$ and using the second order approximation we get $v_{z}=v[1-v_{y}^{2}/(2v^{2})]=v[1-v_{y}^{2}/(2c^{2})]$.
@@
Figure 49: A setup for illustrating the difference between covariant and non-covariant trajectories. The motion of a relativistic electron accelerated by a kicker field. It is assumed for simplicity that $\gamma\gg 1\gg\gamma v_{y}/v$. According to the non-covariant particle tracking, the magnetic field $B\vec{e}_{x}$ is only capable of altering the direction of motion, but not the speed of the electron.
@@
In contrast, covariant particle tracking, which is based on the use of Lorentz coordinates, yields different results for the velocity of the electron. Let us consider a composition of passive Lorentz transformations that track the motion of the relativistic electron accelerated by the kicker field. Let the $S$ be the lab frame of reference and $S^{\prime}$ a comoving frame with velocity $\vec{v}$ relative to $S$. Upstream of the kicker, the particle is at rest in the frame $S^{\prime}$. To have this, we impose that $S^{\prime}$ is connected to $S$ by the Lorentz boost $L(\vec{v})$, with $\vec{v}$ parallel to the $z$ axis, which transforms a given four vector event $X$ in space-time into $X^{\prime}=L(\vec{v})X$. Let us analyze the particle evolution within the $S^{\prime}$ frame. Our particle is at rest and the kicker is running towards it with velocity $-\vec{v}$. The moving magnetic field of the kicker produces an electric field orthogonal to it. When the kicker interacts with the particle in $S^{\prime}$ we thus deal with an electron moving in the combination of perpendicular electric and magnetic fields.
@@
First, we analyze the textbook presentation of a composition of motions. Let $S"$ be a frame fixed with respect to the particle downstream of the kicker. As is known, the non-collinear Lorentz boosts do not commute. In our second-order approximation, we can neglect the difference between the $\gamma v_{y}/c$ and $\gamma_{z}v_{y}/c$, where $\gamma_{z}=1/\sqrt{1-v_{z}^{2}/c^{2}}$. Here $v_{z}=v(1-\theta_{k}^{2}/2)$ and $\theta_{k}=v_{y}/v=v_{y}/c$ in our (ultrarelativistic) case of interest. Therefore we can use a sequence of two commuting non-collinear Lorentz boosts linking $X^{\prime}$ in $S^{\prime}$ to $X^{\prime\prime}$ in $S^{\prime\prime}$ as $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})X^{\prime}=L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e^{\prime}}_{y}v^{\prime}_{y})X^{\prime}$ in order to discuss the beam motion in the frame $S^{\prime}$ after the kick. Here $\vec{e^{\prime}}_{y}$ and $\vec{e^{\prime}}_{z}$ are unit vectors directed, respectively, along the $x^{\prime}$ and $z^{\prime}$ axis. Note that as observed by an observer on $S^{\prime}$, the axes of the frame $S^{\prime\prime}$ are parallel to those of $S^{\prime}$. The relation $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)X$ presents a step-by-step change from $S$ to $S^{\prime}$ and then to $S"$. For the simple case of parallel velocities, the addition law is $L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)=L(\vec{e}_{z}v_{z})$. The resulting Lorentz boost composition can be represented as $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e}_{z}v_{z})X$ and textbooks state that the magnetic field, in the Lorentz coordinatization, is only capable altering the direction of motion, but not the speed of the electron.
@@
According to textbooks, in relativistic kinematics, the composition of the perpendicular velocities is studied analytically from the composition of Lorentz boosts. We have already discussed the law of composition of velocities in Section 15.5. It was shown that velocities addition in the Lorentz coordinatization is regulated by the Wigner rotation. In the ultrarelativistic approximation we find the simple result $v_{tot}=v_{z}=v[1-v_{y}^{2}/(2c^{2})]$, so that a Lorentz transformation leads to a rotation of the particle velocity $v_{z}$ of the angle $v_{y}/c$ (Fig. 50). It becomes clear that the contradiction is related to the Wigner rotation. As we already discussed in the previous sections, the relation $dt^{\prime}=dt/\gamma$ cannot be used when we deal with the relativistic composition of perpendicular velocities in the Lorentz coordinatization.
@@
It should be noted, however, that there is another satisfactory way of covariant electron tracking. Let us now return to our consideration of the motion of a relativistic electron accelerated by the kicker field and let us discuss the observation of an inertial lab observer without passing from one reference frame to another. We now consider the case when the electron in the lab frame is accelerated by the kicker field. In other words, we consider an active boost. In the case of an active boost, we are dealing with the motion of the same physical system, evolving in time and treated from the point of view of the same reference system. The simplest method of synchronization consists of keeping, without changes, the same set of uniformly synchronized clocks used in the case when the electron was upstream the kicker. This synchronization convention preserves simultaneity and is based on the absolute time (or absolute simultaneity) convention. The usual Galileo rule of addition of velocities is used to determine the Galileo boosts tracking an electron, instant after instant, along its motion in a constant magnetic field.
@@
In Section 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization. The overall combination of Galileo transformation and variable changes yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame. Let us analyze the resynchronization process of the lab distant clocks during the acceleration of the electron. This will allow us to demonstrate a direct relation between the decrease of the electron speed after the kick in Lorentz coordinates and the time dilation phenomenon. As we already remarked, the Lorentz coordinate system is only a mental construction: manipulations with nonexisting clocks are only needed for the application of the usual Maxwell’s equations for synchrotron radiation calculations.
@@
Suppose that upstream the kicker we pick a Lorentz coordinates in the lab frame. Then, an instant after entering the magnetic field, the electron velocity changes of the infinitesimal value $d\vec{v}$ along the $y$-axis. At this first step, Eq.(22) allows us to express the differential $d\vec{v}$ through the differential $dt$ in the Lorentz coordinate system assigned upstream of the kicker. If clock synchronization is fixed, this is equivalent to the application of the absolute time convention. To keep Lorentz coordinates in the lab frame, as discussed before, we need to perform a clock resynchronization by introducing an infinitesimal time shift. The simplest case is when the kick angle $\theta_{k}$ is very small, and we evaluate transformations, working only up to the order $(\theta_{k}\gamma)^{2}$. The restriction to this order provides an essential simplicity of calculations in our case of interest for two reasons. First, relativistic correction to compositions of non-collinear velocity increments does not appear in this expansion order, but only in the order $(\gamma\theta_{k})^{3}$. Second, the time dilation appears in the highest order we use. Thus, Eq.(22) allows us to express the small velocity change $\Delta\vec{v}$ after the kick in the initial Lorentz coordinates system, and to perform clock resynchronization only downstream the kicker. Therefore, after the kick we can consider the composition of two Lorentz boosts along the perpendicular $y$ and $z$ directions. The first boost imparts the velocity $v\theta_{k}\vec{e}_{y}$ to the electron along the $y$-axis and the second boost imparts the additional velocity $-(v\theta_{k}^{2}/2)\vec{e}_{z}$ along the $z$ axis, while the restriction to second order assures that the boosts commute.
@@
Figure 50: The motion of a relativistic electron accelerated by the kicker field. To describe the kicker influence on the electron trajectory as Lorentz transformation, one needs to perform a clock resynchronization by introducing a time shift and change the scale of time.
@@
To keep a Lorentz coordinates system in the lab frame after the kick, which is equivalent to describing the kicker influence on the electron trajectory as Lorentz transformation, we need to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of all clocks, from $t$ to $\gamma_{y}t$, with $\gamma_{y}\simeq 1+\theta_{k}^{2}/2$. It is immediately understood that the speed of the electron downstream of the kicker is no longer independent of the electron motion in the magnetic field (Fig. 50). No relativistic correction to the velocity component along the $y$-axis appears in the second order, but a correction of the longitudinal velocity component, changing $v_{z}$ to $v_{z}/\gamma_{y}$ with $v_{z}=v(1-\theta_{k}^{2}/2)$ and $v_{z}/\gamma_{y}=v(1-\theta_{k}^{2})$. It follows that the total electron speed in the lab frame, after clock resynchronization downstream the kicker, decreases from $v$ to $v(1-\theta_{k}^{2}/2)$. The time dilation does not come into the calculation of the velocity increment but appears in the correction of the initial (relativistic) velocity $\vec{v}=v\vec{e}_{z}$. Note that we discuss particle tracking in the limit of a small kick angle $\gamma v_{y}/c\ll 1$. However, even in this simple case and for a single electron we are able to demonstrate the difference between non-covariant and covariant particle trajectories.
18.3 The Motion of an Electron Accelerated by a Kicker in a Bending Magnet